# Random time-shift approximation enables hierarchical Bayesian inference of mechanistic within-host viral dynamics models on large datasets

### v1

@@ 2.3 Viral load measurement model
The data we assume available for each individual is a time-series of the $\log_{10}$ viral load (log-VL) measurements, denoted $\boldsymbol{y}:=(y_{1},\dots,y_{J})$, where $y_{j}:=y_{t_{j}}$ is the log-VL at time $t_{j}$. In practice the $t_{j}$ could be determined in a variety of ways depending on the data-collection process. For a longitudinal study like that run on the NBA cohort [4], daily observations are taken over a long period of time and only the start point of the time-series is effectively random. In a household study, for example, the first time for each individual in a given household is intrinsically linked to both the household being first tested and could come after individuals are infectious, meaning that the first times are random for each index case [37]. Additionally, the timings of the test between individuals are likely to be non-equidistant and also prone to variability as well. The data is right-censored at a threshold $\eta$ (i.e. $y_{j}=\eta$ means that $y_{j}\leq\eta$) which describes the limit of detection of the assay used to measure the viral load [4, 11, 12].
We assume that the data for a single individual are noisy measurements of some underlying VL trajectory. Once more we note that the time-shift, $\tau$, is a random variable that is a function of the VL model parameters, $\boldsymbol{\theta}$ and that $V_{d}(t,\boldsymbol{\theta})$ now has explicit dependence on the VL model parameters $\boldsymbol{\theta}$. Inclusion of the time-shift means that the modelled trajectory is effectively translated in time and so the modelled trajectory is effectively delayed or advanced depending on the sign of the time-shift. In the case where there is a slow start to infection (i.e. a negative time-shift) the modelled trajectory is translated to the right and so is effectively delayed. To handle inconsistencies with the modelled trajectory from applying the time-shift we simply set the VL over this time to the initial condition. For the case where the time-shift is positive, the opposite is true and the modelled trajectory is effectively translated to the left. In this case we can use the modelled trajectory without changes. The only other consideration is that the time $t$ we evaluate the modelled trajectory at must be after the infection time, i.e. $t\geq t_{0}$. This can be summarised by defining
@@
Let $\boldsymbol{z}(\tau):=(z_{1}(\tau),\dots,z_{J}(\tau))$ denote the modelled viral load at the observation times $(t_{1},\dots,t_{J})$ for a particular time-shift $\tau$, that is $z_{j}(\tau):=z_{t_{j}}(\tau):=z(t_{j},\tau,\boldsymbol{\theta})$. Throughout this work we will use the convention that normal distributions are parameterised in terms of mean and standard deviation, i.e. $\mathcal{N}(\mu,\sigma)$, as this parallels the formulation in the code. We assume that the observations are noisy measurements of the modelled trajectory, $y_{j}=z_{j}(\tau)+\epsilon_{j}$ and $\epsilon_{j}\sim\mathcal{N}(0,\kappa)$ for $j=1,\dots,J$ and some standard deviation $\kappa$, which is assumed to be unknown.
@@ 2.4 Mixed effects model
In this section we provide the general form (deferring details of the calculation of the likelihood and specific priors to Sections 2.6 and 2.5, respectively) of our mixed effects model. This includes detailing the data and writing down the posterior distribution.
We assume a simple experimental setup of a sample of $N$ individuals where $\boldsymbol{\theta}_{i}$ denotes the VL model parameters for individual $i$ and $\boldsymbol{y}_{i}=(y_{i,1},\dots,y_{i,J_{i}})$ denotes the time-series of their measured VL, which consists of $J_{i}$ observations. Define $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{N}\}$ and $\mathcal{D}=\{\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{N}\}$ as the collections of parameters and data, respectively. We assume that the subject specific parameters, $\boldsymbol{\theta}_{i}$’s, are distributed according to some population distribution on what we refer to as shared parameters $\boldsymbol{\phi}$ (the exact relationships are defined in Section 2.6). Relationships between the components in the model are depicted in the plate diagram shown in Fig. 2 where the direction of the arrows dictates dependencies between parameters, hyper-parameters and data.
$\boldsymbol{y}_{i}$$\boldsymbol{z}_{i}(t)$$\tau$$\boldsymbol{\theta}_{i}$$\boldsymbol{\phi}$$\kappa$$N$
Fig 2: Plate diagram of the relationship between components in the model. The plate indicates individual level variables that are indexed by the number of individuals, $N$. The parameters outside the plate, $\boldsymbol{\phi}$ and $\kappa$, are shared across all individuals.
The joint posterior distribution of the model parameters and the shared parameters is given by
$f(\boldsymbol{\Theta},\boldsymbol{\phi},\kappa\,|\,\mathcal{D})\propto f(\mathcal{D}\,|\,\boldsymbol{\Theta},\kappa)f(\boldsymbol{\Theta}|\boldsymbol{\phi})f(\boldsymbol{\phi})f(\kappa),$ (7)
where $f(\mathcal{D}\,|\,\boldsymbol{\Theta},\kappa)$ is the joint parameter-likelihood, and $f(\boldsymbol{\Theta}|\boldsymbol{\phi})$, $f(\boldsymbol{\phi})$, and $f(\kappa)$ are prior distributions that are detailed in Section 2.6. Since the individuals are conditionally independent given their unique parameters, $\boldsymbol{\theta}_{i}$, we can factorise the joint-likelihood as
$f(\mathcal{D}\,|\,\boldsymbol{\Theta},\kappa)=\prod_{i=1}^{N}f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa),$ (8)
where $f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)$ is the marginal likelihood function for individual $i$, as it is obtained by marginalising over the random time-shifts (see Section 2.5.1 for details). Conditional on the hyper-parameters, the prior $f(\boldsymbol{\Theta}\,|\,\boldsymbol{\phi})$ can be factorised similarly, which results in a simpler form of the posterior
@@
The difficult part of evaluating Eq. (9) is in the calculation of the marginal likelihood terms, due to the need to marginalise over the distribution of trajectories as seen in Fig. 2.
@@ 2.5 Likelihood approximations
2.5.1 Marginal likelihood
The marginal likelihood (likelihood for individual $i$’s data, $\boldsymbol{y}_{i}$, given the parameters, $\boldsymbol{\theta}_{i}$) is given by $f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)$. Since the observation process depends on an unobserved random time-shift $\tau$, we marginalise over $\tau$
[…]
The likelihood given $\tau$ depends on the system’s trajectory $\boldsymbol{z}_{i}(\tau)$, which is deterministic for a given $\tau$ and $\boldsymbol{\theta}$, so we can rewrite the conditional likelihood as
$f(\boldsymbol{y}_{i}\,|\,\tau,\boldsymbol{\theta}_{i},\kappa)=f(\boldsymbol{y}_{i}\,|\,\boldsymbol{z}_{i}(\tau),\boldsymbol{\theta}_{i},\kappa).$ (11)
The distribution of the random time-shift $F(\tau\,|\,\boldsymbol{\theta}_{i})$ has a single point mass at $\tau=-\infty$, of size $q^{\star}(\boldsymbol{\theta}_{i})$, corresponding to extinction, where the likelihood of observing any data is zero. Excluding this case, the remaining portion of $F(\tau\,|\,\boldsymbol{\theta}_{i})$ is absolutely continuous with density $f(\tau\,|\,\boldsymbol{\theta}_{i})$, allowing us to write
[…]
Observations $y_{i,j}$ correspond to viral load (VL) measurements at times $t_{j}$, which are assumed to be independent given the model trajectory. The deterministic trajectory $z_{i,j}(\tau):=z(t_{j},\tau,\boldsymbol{\theta}_{i})$ represents the modelled VL at time $t_{j}$ for individual $i$. Given $z_{i,j}(\tau)$, the conditional likelihood of the observed data $\boldsymbol{y}_{i}$ is
@@ 2.5.1 Marginal likelihood
where $\mathcal{N}_{\textrm{PDF}}(y_{i,j}\,|\,z_{i,j}(\tau),\boldsymbol{\theta},\kappa)$ and $\mathcal{N}_{\textrm{CDF}}(\eta\,|\,z_{i,j}(\tau),\boldsymbol{\theta},\kappa)$ are the probability density function, and cumulative distribution function, respectively, of a $\mathcal{N}(z_{i,j}(\tau),\kappa)$ random variable. The CDF terms account for observations below the detection threshold [11]. To distinguish this from other likelihood formulations, we refer to Eq. (13) as the path-likelihood for a single individual. The marginal likelihood, Eq. (12), can be evaluated by substituting in the path-likelihood, Eq. (13), but this is inefficient due to the numerical integration of an interpolated ODE solution. In Section 2.5.2 we develop a faster approximation which replaces this calculation.
@@
We note here that if the VL model is replaced with the deterministic model without the time-shift, then this is equivalent to the time-shift density collapsing to a point mass at $\tau=0$ and the removal of the extinction probability factor. Hence, the likelihood is simply $f(\boldsymbol{y}_{i}\,|\,\boldsymbol{z}_{i}(0),\boldsymbol{\theta}_{i},\kappa)$ as often seen in VL kinetics modelling [8, 5, 11, 38].
2.5.2 Laplace approximation of the path likelihood
The marginal likelihoods (Eq. (12)) do not admit a simple closed form expression and so we resort to numerical integration. We use an adaptive Gauss–Kronrod quadrature for numerical integration [39] as it is highly efficient and demonstrated strong performance in our testing. The challenge with evaluating the marginal likelihood directly is that it relies on repeated evaluation of the modelled viral-load. Repeating this calculation is necessary as we are integrating over $\tau$ which appears inside the ODE solution (i.e. $\boldsymbol{z}_{i}(\tau)$). To evaluate this, the ODE solution is interpolated and calling this inside a numerical integrator is computationally expensive.
To address this, we consider a Laplace approximation of the interior of the integral of the marginal likelihood [40, Chapter 27]. We evaluated two approaches: (1) constructing a Laplace approximation for the path-likelihood alone and then numerically integrating, or (2) constructing a Laplace approximation for the product of the path-likelihood and the time-shift distribution, and deriving a closed-form expression. The approximation of the path-likelihood alone was initially considered the more appropriate approach, as the path-likelihood is the product of normal PDFs (and few CDF values), and therefore, it is approximately a scaled Normal distribution (note that the scaling means this is a function, not a density). In contrast, the time-shift distribution is skewed (see Fig. 1). However, in our testing, we found that these two approaches produced comparable results, with the Laplace approximation of the product of the path-likelihood and time-shift distribution reducing the evaluation time of the likelihood by approximately 30%.
We begin the construction of the Laplace approximation by noting that the data, $\boldsymbol{y}_{i}$, VL model parameters, $\boldsymbol{\theta}_{i}$, and $\kappa$ are assumed to be known and fixed. Define the functions
$g(\tau)=f(\boldsymbol{y}_{i}\mid\boldsymbol{z}_{i}(\tau),\boldsymbol{\theta}_{i},\kappa)f(\tau\mid\boldsymbol{\theta}_{i}),\quad h(\tau)=\log g(\tau).$
Suppose there is a global maximum
$\tau_{0}=\arg\max_{\tau}\,g(\tau)=\arg\max_{\tau}\,h(\tau),$
and define
$H(\tau)=\sqrt{-\frac{1}{h^{\prime\prime}(\tau)}},$
noting that the function is assumed to be concave (implicit in applying the Laplace approximation), which implies $h^{\prime\prime}(\tau)<0$ around $\tau_{0}$, and hence $H(\tau)>0$. Taylor expanding $h(\tau)$ about $\tau_{0}$ and noting that $h^{\prime}(\tau_{0})=0$ (since $\tau_{0}$ is the maximum), we get
$h(\tau)\approx h(\tau_{0})+\frac{1}{2}h^{\prime\prime}(\tau_{0})(\tau-\tau_{0})^{2}=h(\tau_{0})-\frac{1}{2H(\tau_{0})^{2}}(\tau-\tau_{0})^{2}.$
Then, we have
$g(\tau)=\exp{(h(\tau))}\approx g(\tau_{0})\exp{\left(-\frac{(\tau-\tau_{0})^{2}}{2H(\tau_{0})^{2}}\right)}.$
Since $H(\tau_{0})>0$, this is the kernel of a $\mathcal{N}(\tau_{0},H(\tau_{0}))$ random variable scaled by $g(\tau_{0})=f(\boldsymbol{y}_{i}\mid\boldsymbol{z}_{i}(\tau_{0}),\boldsymbol{\theta}_{i},\kappa)$. Substituting into Eq. (12), the Laplace approximation is
[…]
Since the interior of the integrand is a Normal kernel, a closed-form approximation is given by
$f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)\approx\sqrt{2\pi}(1-q^{\star}(\boldsymbol{\theta}_{i}))f(\boldsymbol{y}_{i}\mid\boldsymbol{z}_{i}(\tau_{0}),\boldsymbol{\theta}_{i},\kappa)f(\tau_{0}\mid\boldsymbol{\theta}_{i})H(\tau_{0}).$ (14)
The mean of the Gaussian, $\tau_{0}$, can be estimated using numerical optimisation. The scale term $H(\tau_{0})$ has a more complicated form but can be computed through either numerical differentiation or automatic differentiation. In this work, all our code is written in the Julia programming language [41], and the package ForwardDiff.jl [42] enables efficient computation of exact derivatives through forward mode automatic differentiation for such codes.
@@ 2.5.2 Laplace approximation of the path likelihood
In Section 4 of S1 text, we plot profile likelihoods—i.e., the likelihood with all parameters fixed except for one, which is allowed to vary—for a simulated VL trajectory to show the accuracy of the approximation in Eq. (14) versus the exact calculation in Eq. (12). The profile likelihoods had a maximum (absolute) error of $<10^{-7}$ at the tested points, which indicates strong agreement between our Laplace approximation and the exact likelihood.
@@ 2.6 Priors
In this section, we detail our choices of priors and fixed parameters. To ensure broad applicability, we center our priors at values consistent with studies on other viral infections (e.g., influenza) and COVID-19. To account for uncertainty in these estimates, we use weakly informative priors with higher variance. Prior predictive checks confirm that the effective priors on individual parameters remain weakly informative. It is important to note that our primary focus is methodological—demonstrating how to incorporate stochasticity into WHVD models—rather than obtaining precise parameter estimates. Additionally, because our hierarchical model is fit to a larger number of individuals than in previous studies, the population-level estimate is better informed by the data, making it less reliant on the prior. While hierarchical models can sometimes be more sensitive to prior misspecification, the larger sample size in this case helps mitigate this issue as the data is more informative.
Due to identifiability concerns, some VL parameters are fixed at values that are well established in the literature. We assume that the eclipse phase of newly infected cells, $k_{i}$, and the clearance rate of virus, $c_{i}$, are fixed at $k_{i}=4\textrm{ d}^{-1}$ and $c_{i}=10\textrm{ d}^{-1}$ for $i=1,\dots,N$, respectively [5]. As outlined in Section 2.2, we parameterise the model in terms of the basic reproduction number, $R_{0}$. Table 2 gives some estimates of $R_{0}$ for COVID-19 and Influenza as well as the dataset size (number of individuals) from the literature. None of these studies implemented hierarchical models and instead fit each individual time-series.
@@
$\mu_{\delta}\sim\textrm{Gamma}(26,0.05)$
which yields a mean rate of removal of $1.3~{}d^{-1}$ with a 95% quantile interval of $(0.85,1.85)$.
The parameter $\mu_{\rho}$ is difficult to assign a prior, as it has identifiability issues similar to $\beta$ [5, 18]. To address this, we assume
@@
which has median $3.5~{}d^{-1}$ (roughly based on [5]) and a long upper tail with a 95% quantile interval of $(1.1,8.77)$. This distribution is deliberately weakly informative and reflects values estimated for COVID-19 in other studies [45, 43].
@@
The hyper-prior for $\sigma_{R_{0}}$ and $\sigma_{\rho}$ have a larger standard deviation compared with $\sigma_{\delta}$ as we found in our testing that a $\mathcal{N}_{+}(0,1)$ over-constrained the fitted estimates which caused excessive shrinkage of the individual level parameter estimates to the population mean.
We use a non-centred parameterisation (NCP) for the individual-level parameters to enhance posterior sampling efficiency [46]. Specifically, the individual-specific parameters $R_{0_{i}},\delta_{i},$ and $\rho_{i}$ are modelled as
$R_{0_{i}}=\mu_{R_{0}}+\sigma_{R_{0}}z_{R_{0_{i}}},\quad\delta_{i}=\mu_{\delta}+\sigma_{\delta}z_{\delta_{i}},\quad\rho_{i}=\mu_{\rho}+\sigma_{\rho}z_{\rho_{i}}$
where $z_{R_{0_{i}}},z_{\delta_{i}},z_{\rho_{i}}\sim\mathcal{N}(0,1)$. This parameterisation facilitated more efficient posterior sampling.
The scale parameter of the measurement noise, $\kappa$, is also assigned a weakly informative prior
@@
This choice results in a median time of infection (pre-peak VL) of $-5.9$ days. This prior assumes infection occurs most likely before peak VL but allows for post-peak infection times, which may occur in some cases due to measurement variability or delayed detection. Typical scenarios where this prior is useful is when the peak VL is a low measurement or all measurements are tightly clustered in time so it’s difficult to distinguish whether the peak is early or late on in the infection. This distribution is consistent with the incubation periods of COVID-19 (pre-Omicron variant) [47] when assuming that peak infectiousness (i.e. peak VL) corresponds to symptom onset. For ease of notation we will denote the set of all hyper-parameters as $\boldsymbol{\phi}=(\mu_{R_{0}},\mu_{\delta},\mu_{\rho},\sigma_{R_{0}},\sigma_{\delta},\sigma_{\rho})$.

### v2

@@ 2.3 Viral load measurement model
The data we assume available for each individual is a time-series of the $\log_{10}$ viral load (log-VL) measurements, denoted $\boldsymbol{y}\mathrel{\mathop{\ordinarycolon}}=(y_{1},\dots,y_{J})$, where $y_{j}\mathrel{\mathop{\ordinarycolon}}=y_{t_{j}}$ is the log-VL at time $t_{j}$. In practice the $t_{j}$ could be determined in a variety of ways depending on the data-collection process. For a longitudinal study like that run on the NBA cohort [4], daily observations are taken over a long period of time and only the start point of the time-series is effectively random. In a household study, for example, the first time for each individual in a given household is intrinsically linked to both the household being first tested and could come after individuals are infectious, meaning that the first times are random for each index case [39]. Additionally, the timings of the test between individuals are likely to be non-equidistant and also prone to variability as well. The data is left-censored at a threshold $\eta$ (i.e. $y_{j}=\eta$ means that $y_{j}\leq\eta$) which describes the limit of detection of the assay used to measure the viral load [4, 11, 12].
We assume that the data for a single individual are noisy measurements of some underlying VL trajectory. Once more we note that the time-shift, $\tau$, is a random variable that is a function of the VL model parameters, $\boldsymbol{\theta}$ and that $V_{d}(t,\boldsymbol{\theta})$ now has explicit dependence on the VL model parameters $\boldsymbol{\theta}$. Inclusion of the time-shift means that the modelled trajectory is effectively translated in time and so the modelled trajectory is delayed or advanced depending on the sign of the time-shift. In the case where there is a slow start to infection (i.e. a negative time-shift) the modelled trajectory is translated to the right and so is effectively delayed. To handle inconsistencies with the modelled trajectory from applying the time-shift we simply set the VL over this time to the initial condition. For the case where the time-shift is positive, the opposite is true and the modelled trajectory is effectively translated to the left. In this case we can use the modelled trajectory without changes. The only other consideration is that the time $t$ we evaluate the modelled trajectory at must be after the infection time, i.e. $t\geq t_{0}$. This can be summarised by defining
@@
Let $\boldsymbol{z}(\tau)\mathrel{\mathop{\ordinarycolon}}=(z_{1}(\tau),\dots,z_{J}(\tau))$ denote the modelled viral load at the observation times $(t_{1},\dots,t_{J})$ for a particular time-shift $\tau$, that is $z_{j}(\tau)\mathrel{\mathop{\ordinarycolon}}=z_{t_{j}}(\tau)\mathrel{\mathop{\ordinarycolon}}=z(t_{j},\tau,\boldsymbol{\theta})$. Throughout this work we will use the convention that normal distributions are parameterised in terms of mean and standard deviation, i.e. $\mathcal{N}(\mu,\sigma)$, as this parallels the formulation in the code. We assume that the observations are noisy measurements of the modelled trajectory, $y_{j}=z_{j}(\tau)+\epsilon_{j}$ and $\epsilon_{j}\sim\mathcal{N}(0,\kappa)$ for $j=1,\dots,J$ and some standard deviation $\kappa$, which is assumed to be unknown.
@@ 2.4 Mixed effects model
In this section we provide the general form (deferring details for priors and the calculation of the likelihood to Sections 2.5 and 2.6, respectively) of our mixed effects model. This includes detailing the data and writing down the posterior distribution.
We assume a simple experimental setup of a sample of $N$ individuals where $\boldsymbol{\theta}_{i}$ denotes the VL model parameters for individual $i$ and $\boldsymbol{y}_{i}=(y_{i,1},\dots,y_{i,J_{i}})$ denotes the time-series of their measured VL, which consists of $J_{i}$ observations. Define $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{N}\}$ and $\mathcal{D}=\{\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{N}\}$ as the collections of parameters and data, respectively. We assume that the subject-specific parameters, $\boldsymbol{\theta}_{i}$, are drawn from a population distribution governed by what we refer to as the shared parameters $\boldsymbol{\phi}$ (the exact relationships are defined in Section 2.6). The joint posterior is given by,
@@
The difficult part of evaluating Eq. (5) is in the calculation of the marginal likelihood terms, $f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)$, due to the need to marginalise over the distribution of trajectories.
@@ 2.5 Likelihood approximations
The marginal likelihood (likelihood for individual $i$’s data, $\boldsymbol{y}_{i}$, given the parameters, $\boldsymbol{\theta}_{i}$) is given by $f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)$. Since the observation process depends on an unobserved random time-shift $\tau$, we marginalise over $\tau$, considering only the non-extinction case
$f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)=\int f(\boldsymbol{y}_{i}\,|\,\tau,\boldsymbol{\theta}_{i},\kappa)f(\tau\,|\,\boldsymbol{\theta}_{i})\operatorname{d\!}{\tau}.$ (6)
The conditional likelihood given $\tau$ depends on the system’s trajectory $\boldsymbol{z}_{i}(\tau)$, which is deterministic for a given $\tau$ and $\boldsymbol{\theta}$, so we can rewrite the marginal likelihood as
$f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)=\int f(\boldsymbol{y}_{i}\,|\,\boldsymbol{z}_{i}(\tau),\boldsymbol{\theta}_{i},\kappa)f(\tau\,|\,\boldsymbol{\theta}_{i})\operatorname{d\!}{\tau}.$ (7)
Observations $y_{i,j}$ correspond to viral load (VL) measurements at times $t_{j}$, which are assumed to be independent given the model trajectory. The deterministic trajectory $z_{i,j}(\tau)\mathrel{\mathop{\ordinarycolon}}=z(t_{j},\tau,\boldsymbol{\theta}_{i})$ represents the modelled VL at time $t_{j}$ for individual $i$. Given $z_{i,j}(\tau)$, the conditional likelihood of the observed data $\boldsymbol{y}_{i}$ is
@@
where $\mathcal{N}_{\textrm{PDF}}(y_{i,j}\,|\,z_{i,j}(\tau),\boldsymbol{\theta},\kappa)$ and $\mathcal{N}_{\textrm{CDF}}(\eta\,|\,z_{i,j}(\tau),\boldsymbol{\theta},\kappa)$ are the probability density function, and cumulative distribution function (CDF), respectively, of a $\mathcal{N}(z_{i,j}(\tau),\kappa)$ random variable. The CDF terms account for observations below the detection threshold [11]. To distinguish this from other likelihood terms, we refer to Eq. (8) as the path-likelihood for a single individual. We note here that if the VL model is replaced with the deterministic model without the time-shift, then this is equivalent to the time-shift density collapsing to a point mass at $\tau=0$. Hence, the likelihood is simply $f(\boldsymbol{y}_{i}\,|\,\boldsymbol{z}_{i}(0),\boldsymbol{\theta}_{i},\kappa)$ as often seen in VL kinetics modelling [8, 5, 11, 40].
The marginal likelihood, Eq. (7), can be evaluated by substituting the path-likelihood, Eq. (8), and integrating. However, it involves an integral over a function that does not admit an analytical solution, so numerical methods are required. Numerical integration requires repeated computation of the modelled viral load as we integrate over $\tau$, which appears inside the ODE solution (i.e., $\boldsymbol{z}_{i}(\tau)$). Evaluating the ODE at each integration step, even with interpolation, is computationally expensive.
To address this, we consider a Laplace approximation [41, Chapter 27] of the interior of the integral of the marginal likelihood. The approximation is constructed by noting that for fixed $\boldsymbol{\theta}_{i}$, $\kappa$ and data $\boldsymbol{y}_{i}$, the marginal likelihood is simply a function of $\tau$. Let
$h(\tau)=f(\boldsymbol{y}_{i}\,|\,\boldsymbol{z}_{i}(\tau),\boldsymbol{\theta}_{i},\kappa)f(\tau\,|\,\boldsymbol{\theta}_{i}).$ (9)
then the Laplace approximation to the marginal likelihood (Eq. (7)) is given by (see Section 4 of S1 text. for derivation)
$f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)\approx\frac{\sqrt{2\pi}h(\tau_{0})}{\sqrt{-h^{\prime\prime}(\tau_{0})}},$ (10)
@@
noting that $h(\tau)$ is assumed to be concave (implicit in applying the Laplace approximation), which implies $h^{\prime\prime}(\tau)<0$ around $\tau_{0}$. The quantity $\tau_{0}$, is the maximiser of Eq. (9) and can be estimated efficiently using optimisation. The denominator, $((-h^{\prime\prime}(\tau_{0}))^{-1/2}$, is the scale term and has a more complicated form but can also be computed through either numerical differentiation or automatic differentiation of Eq. (9). In this work, all our code is written in the Julia programming language [42], and the package ForwardDiff.jl [43] enables efficient computation of exact derivatives through forward mode automatic differentiation. In Section 4 of S1 text., we plot profile likelihoods—i.e., the likelihood with all parameters fixed except for one, which is allowed to vary—for a simulated VL trajectory to show the accuracy of the approximation in Eq. (10) versus the exact calculation in Eq. (7). The profile likelihoods had a maximum (absolute) error of $10^{-7}$ at the tested points, which indicates strong agreement between our Laplace approximation and the exact likelihood.
@@ 2.6 Priors
In this section, we detail our choices of priors and fixed parameters. We center our priors at values consistent with studies of other viral infections (e.g., influenza) and COVID-19, and use weakly informative priors with higher variance to account for uncertainty in these estimates. Prior predictive checks confirm that the effective priors on individual parameters remain weakly informative. Our primary focus in this work is methodological—demonstrating how to incorporate stochasticity into WHVD models and speed up the inference—rather than conducting a thorough analysis of this data. Moreover, because our hierarchical model is fit to a larger number of individuals than in previous studies, the population-level estimates are better informed by the data and less reliant on the prior. While hierarchical models can be sensitive to prior misspecification, the larger sample size here mitigates this concern, as the data dominate the inference.
Due to identifiability concerns, some VL parameters are fixed at values that are well established in the literature. We assume that the eclipse phase of newly infected cells, $k_{i}$, and the clearance rate of virus, $c_{i}$, are fixed at $k_{i}=4\textrm{ d}^{-1}$ and $c_{i}=10\textrm{ d}^{-1}$ for $i=1,\dots,N$, respectively [5]. The individual-level parameters, $R_{0,i}$, $\delta_{i}$, and $\rho_{i}$, are all assumed to be normally distributed (truncated to be non-negative) with unknown mean and scale parameters. Here we detail the hyper-priors on these parameters. As outlined in Section 2.2, we parameterise the model in terms of the basic reproduction number, $R_{0}$. Table 2 gives some estimates of $R_{0}$ for COVID-19 and Influenza as well as the dataset size (number of individuals) from the literature. None of these studies implemented hierarchical models and instead fit each individual time-series.
@@
$\mu_{\delta}\sim\textrm{Gamma}(26,0.05),$
which yields a mean rate of removal of $1.3~d^{-1}$ with a 95% quantile interval of $(0.85,1.85)$. The parameter $\mu_{\rho}$ is difficult to assign a prior, as it has identifiability issues similar to $\beta$ [5, 18]. To address this, we assume
@@
which has median $2.9~d^{-1}$ (roughly based on [5]) and a 95% quantile interval of $(1.44,5.13)$. This distribution is deliberately weakly informative and reflects values estimated for COVID-19 in other studies [46, 44].
@@
The hyper-prior for $\sigma_{R_{0}}$ and $\sigma_{\rho}$ have a larger standard deviation compared with $\sigma_{\delta}$ as we found in our testing that a $\mathcal{N}_{+}(0,1)$ over-constrained the fitted estimates which caused excessive shrinkage of the individual level parameter estimates to the population mean. The standard deviation of the measurement noise, $\kappa$, is also assigned a weakly informative prior
@@
This choice results in a median time of infection (pre-peak VL) of 5.9-5.9 days. This prior assumes infection occurs most likely before peak VL but allows for post-peak infection times, which may occur in some cases due to measurement variability or delayed detection. Typical scenarios where this prior is useful is when the peak VL is a low measurement or all measurements are tightly clustered in time so it’s difficult to distinguish whether the peak is early or late on in the infection. This distribution is consistent with the incubation periods of COVID-19 (pre-Omicron variant) [47] when assuming that peak infectiousness (i.e. peak VL) corresponds to symptom onset. For ease of notation we will denote the set of all hyper-parameters as
$\boldsymbol{\phi}=(\mu_{R_{0}},\mu_{\delta},\mu_{\rho},\sigma_{R_{0}},\sigma_{\delta},\sigma_{\rho}).$